# Rethinking Spectral Graph Neural Networks with Spatially Adaptive Filtering

### v3

@@ 5.2. Non-local Spatial Aggregation
where $\eta$ refers to the update rate and $\mathbf{Z}^{(0)}=f_{\varphi}(\mathbf{X})$. The iteratively aggregated results are denoted as $\mathbf{Z}_{a}$. Recognizing the potential noise from the non-local nature of $\hat{\mathbf{A}}^{\text{new}}$, we apply a sparsification technique, leveraging a positive threshold $\epsilon$, and retain only essential elements outside the $\left[-\epsilon,\epsilon\right]$ interval. For clarity, this refined model is referred to as SAF-$\epsilon$.
Figure 3. Illustration of the proposed SAF framework, where varying node colors represent different node labels.
5.3. Node-wise Prediction Amalgamation
@@
To leverage information from different graph domains, we employ an attention mechanism, allowing nodes to determine the importance of each space. This mechanism produces pairwise weights for a nuanced amalgamation during prediction. Specifically, the weight pair is computed as $\boldsymbol{\kappa}_{f}=\text{Sigmoid}(\mathcal{P}_{f}(\mathbf{Z}_{f}))$, $\boldsymbol{\kappa}_{a}=\text{Sigmoid}(\mathcal{P}_{a}(\mathbf{Z}_{a}))$ where $\boldsymbol{\kappa}_{f},\boldsymbol{\kappa}_{a}\in\mathbb{R}^{N}$ contain the weights for each node, and $\mathcal{P}_{f}(\cdot)$ and $\mathcal{P}_{a}(\cdot)$ are two different mappings from $\mathbb{R}^{C}$ to $\mathbb{R}$. For simplicity, we implement them using two one-layer MLPs. Given domain predictions $\mathbf{Y}_{f},\mathbf{Y}_{a}\in\mathbb{R}^{C}$, the final model prediction is attained as
(7) $\mathbf{Y}=\mathrm{diag}(\boldsymbol{\kappa}_{f})\cdot\mathbf{Y}_{f}+\mathrm{diag}(\boldsymbol{\kappa}_{a})\cdot\mathbf{Y}_{a}$
where a normalization $\left[\boldsymbol{\kappa}_{f},\boldsymbol{\kappa}_{a}\right]\leftarrow\frac{\left[\boldsymbol{\kappa}_{f},\boldsymbol{\kappa}_{a}\right]}{\max\{\|\{\boldsymbol{\kappa}_{f},\boldsymbol{\kappa}_{a}\}\|_{1},\delta\}}$ is performed beforehand to maintain $\boldsymbol{\kappa}_{f}+\boldsymbol{\kappa}_{a}=\mathbf{1}$ with small value $\delta$ preventing zero division. Similar schemes can be founded in works (Sagi and Rokach, 2018; Wang et al., 2020; Zhu et al., 2020a).
@@
Table 1. Semi-supervised node classification accuracy (%) $\pm$ 95% confidence interval.
@@
Improv.333 Improv. indicates the relative improvement of SAF over its base model, BernNet (He et al., 2021). For alternative implementation using ChebNetII (He et al., 2022), please refer to Appendix E.4. 14.56% 9.68% 15.37% 13.99% 3.97% 1.62% 1.86% 0.68%
Table 2. Full-supervised node classification accuracy (%) $\pm$ 95% confidence interval.
@@
Improv.333 Improv. indicates the relative improvement of SAF over its base model, BernNet (He et al., 2021). For alternative implementation using ChebNetII (He et al., 2022), please refer to Appendix E.4. 6.77% 12.61% 2.13% 1.15% 1.27% 1.36% 1.37% 3.01%
5.4. Complexity Analysis
SAF augments spectral GNNs with non-local aggregation and node-wise amalgamation. The first part entails creating a new graph and information propagation. In SAF-$\epsilon$, these two steps are separated by sparsification, culminating in $\mathcal{O}(N^{3}+N^{2}+nnz(\hat{\mathbf{A}}^{\text{new}})d)$ complexity, where $nnz$ denotes non-zero element count. Conversely, SAF, viewing non-local aggregation holistically, can reduce complexity to $\mathcal{O}(2dN^{2}+dN)$ when $d\ll N$. For node-wise amalgamation, its parallelizable nature ensures computational efficiency. Our method also requires a precomputation of eigendecomposition [^2] For more on its modern applications and discussions, please refer to Appendix B (including Specformer (Bo et al., 2023) & FE-GNN (Sun et al., 2023) featured in our experiments) and E.5. which, naively complex at $\mathcal{O}(N^{3})$, can be reduced to $\mathcal{O}(m^{2}+nnz(\hat{\mathbf{L}})m)$ using Lanczos method (Lanczos, 1950) for larger graphs with $m\ll N$ denoting iterative steps and is reusable for both training and inference. We present empirical studies on both time and space overheads in Appendix E.5.
6. Experiments
6.1. Datasets and Experimental Setup
6.1.1. Datasets
We evaluate models over 13 real-world datasets from various domains. These include three well-known homophilic graphs: Cora, Citeseer, and Pubmed (Sen et al., 2008), five commonly used heterophilic graphs: Chameleon, Squirrel (Rozemberczki et al., 2021), Cornell, Texas (Pei et al., 2020), and Actor (Tang et al., 2009), as well as five recently introduced benchmarks: Minesweeper (synthetic graph), Tolokers (crowdsourcing platform worker network), Amazon-ratings (product co-purchasing), Roman-empire (word dependency graph) (Platonov et al., 2023b), and Penn94 (Lim et al., 2021) (social network). Detailed statistics are provided in Appendix C.
6.1.2. Baselines
We compare SAF with 22 models: (1) MLP; (2) Basic GNNs: GCN (Kipf and Welling, 2017) & APPNP (Gasteiger et al., 2019); (3) Spectral GNNs: ARMA (Bianchi et al., 2021), GPR-GNN (Chien et al., 2021), BernNet (He et al., 2021), ChebNetII (He et al., 2022), JacobiConv (Wang and Zhang, 2022), Specformer (Bo et al., 2023), LON-GNN (Tao et al., 2023) & OptBasisGNN (Guo and Wei, 2023b); (4) Spatial GNNs: GCNII (Chen et al., 2020), PDE-GCN (Eliasof et al., 2021), GEN (Wang et al., 2021), NodeFormer (Wu et al., 2022), GloGNN++ (Li et al., 2022) & MGNN (Cui and Wei, 2023); (5) Unified GNNs: GNN-LF (Zhu et al., 2021), GNN-HF (Zhu et al., 2021), ADA-UGNN (Ma et al., 2021), FE-GNN (Sun et al., 2023) & ClenshawGCN (Guo and Wei, 2023a).
6.1.3. Setup
To follow (He et al., 2021, 2022; Wang and Zhang, 2022), we fix $K=10$. For each dataset, we perform a grid search to tune the hyper-parameters of all models. With the best hyper-parameters, we train models with Adam optimizer (Kingma and Ba, 2014) in 1,000 epochs using early-stopping strategy and a patience of 200 epochs, and report the mean classification accuracies with a 95% confidence interval on 10 random data splits. More experimental details can be found in Appendix D, and the codes will be made available if the paper could be accepted.
6.2. Overall Evaluation
6.2.1. Semi-supervised Node Classification.
@@ 6.2. Overall Evaluation
In this task, we follow the experimental protocol established by (He et al., 2022) and compare our models with MLP, two basic GNNs, eight popular polynomial spectral GNNs, and five unified GNNs. For data splitting on homophilic graphs (Cora, Citeseer, and Pubmed), we apply the standard division (Yang et al., 2016) with 20 nodes per class for training, 500 nodes for validation, and 1,000 nodes for testing. On the other five heterophilic graphs, we leverage the sparse splitting (Chien et al., 2021) with 2.5%/2.5%/95% samples respectively for training/validation/testing. The results are reported in Table 1, where the best results are bold and the underlined letters denote the second highest accuracy. We first observe that both SAF and SAF-$\epsilon$ substantially boosts its base model, BernNet, with gains reaching a notable 15.37%. This impressive enhancement is credited to their capacity to effectively exploit the task-beneficial information, which is implicitly encoded by spectral filtering in the spatial domain. This ability is particularly advantageous in contexts with limited supervision, where it allows effective leveraging of extra prior knowledge during training. Generally, our models outperform competitors on all datasets except for Chameleon and Squirrel, where SAF maintains a second-place rank with considerable improvements on BernNet by 14.56% and 9.68%. In these cases, ChebNetII initially surpasses our model, yet, with more training samples, our SAF manages to beats it by margins of 3.93% and 6.28% (see Table 2). Moreover, it can be seen that SAF-$\epsilon$ averagely delivers better results than SAF, benefiting from its thresholding sparsity that reduces non-local noise for efficient graph learning. However, this enhancement also incurs higher computational costs, as illustrated in both Section 5 and Appendix E.5.
6.2.2. Full-supervised Node Classification
444 We borrow this terminology from (Chen et al., 2020; He et al., 2022) to denote 60%/20%/20% data splitting.
@@ 6.2.2. Full-supervised Node Classification
To bolster our evaluation, we expand the previously compared baselines to include six cutting-edge spatial GNN models: GCNII & PDE-GCN for capturing long-range dependency, GEN & MGNN for handling graph heterophily, and NodeFormer & GloGNN++ addressing both. For all datasets, we randomly divide them into 60%/20%/20% for training/validation/testing by following (He et al., 2021, 2022). Table 2 summarizes the mean classification accuracies. Our methods demonstrate superior performance across most datasets, with an exception on Squirrel where they achieve comparable results to Specformer. This notable performance is primarily attributed to our SAF’s effective non-local aggregation, utilizing signed edge weights to model global label relationships. This enables our methods to outperform GNNs that are specifically tailored for long-range dependency and/or graph heterophily. Besides the eight standard datasets for node classification, our study also extends to five new benchmarks (Lim et al., 2021; Platonov et al., 2023b) that focus on graph heterophily. Due to space limit, we present detailed results in Appendix E.3.
Figure 4. Ablation study of SAF framework on six datasets. Refer to Table 8 in the Appendix for comprehensive results.
6.3. Ablation Study
This section aims to validate our designs by comparing SAF with its three ablated variants – SAF w/o Atte., SAF w/o Spec., and SAF w/o Spat. – in full-supervised node classification. Specifically, Atte., Spec., and Spat. respectively refers to: attention mechanism in “Node-wise Prediction Amalgamation”, “Non-negative Spectral Filtering”, and “Non-local Spatial Aggregation”. For SAF w/o Atte., we remove the attention mechanism and equally blend predictions from different domains. SAF w/o Spec. abandons the spectral filtering phase, practically setting $\boldsymbol{\kappa_{f}}=\mathbf{0},\boldsymbol{\kappa_{a}}=\mathbf{1}$ in Eq. (7). As the SAF w/o Spat. configuration is equivalent to BernNet model, the corresponding results are posted directly. From Figure 4, we can draw several insights: 1) The impact of Atte. module on our SAF varies by datasets, e.g., on Chameleon and Squirrel, showing a slight performance reduction upon its removal. This observation aligns with our observation that their optimal attention values are close to an even split, as suggested in Figures 5(a) and 9(a) Conversely, Cora dataset exhibits a notable drop, due to its optimal attention weights being far from even, as depicted in Figure 5(b). 2) Spectral filtering (Spec. module) remains vital for discriminative node representation learning. Specifically, the quality of the adapted new graph fundamentally hinges on the graph spectral filters’ training, as underscored by their theoretical interaction in Eq. (6). Practically, the absence of spectral filtering markedly reduces model accuracy, confirming its importance in SAF. 3) This visualization not only reaffirms the pivotal role of the non-local aggregation (Spat. module), but also underscores its position as the most crucial component in advancing spectral GNNs within the SAF framework.
6.4. Analysis of Attention Trends
We analyze the changing trends of the pair-wise attention weights during training SAF on Squirrel and Cora datasets. From Figure 5, the average weights for filtering and aggregation start similarly but diverge throughout training, showing different trends in heterophilic and homophilic graphs. On the heterophilic graph Squirrel, both weights converge to similar values, demonstrating their mutual importance in modeling complex connectivity. Conversely, $\boldsymbol{\kappa}_{f}$ becomes dominant on the homophilic graph Cora due to the sufficiency of node proximity information for label prediction, thereby diminishing the relevance of $\boldsymbol{\kappa}_{a}$ and non-local aggregation.
6.5. Parameter Study
This section presents the sensitivity analysis of hyper-parameters including $\tau$, $\eta$, $\epsilon$, and $L$. Figure 6 visualizes how varying these parameters within a broad range influences learning performance, showcasing our model’s robust stability over diverse settings. Beyond empirical observation, we also provides deeper insights into parameter understanding and rationalizes the chosen ranges for parameter searching: 1) The scaling parameter $\tau=\frac{\alpha}{1-\alpha}$, crucial in new graph construction in Eq. (6), stems from the trade-off parameter $\alpha\in(0,1)$ within the graph optimization problem in Eq.(3). While theoretically we have $0=\frac{0}{1-0}<\tau<\frac{1}{1-1}=\infty$, practical considerations for extracting structural information suggest a larger penalty on the trace objective term $\mathrm{tr}(\mathbf{Z}^{T}\gamma_{\theta}(\hat{\mathbf{L}})\mathbf{Z})$, i.e., keeping $\alpha<0.5$, thereby limiting $\tau<\frac{0.5}{1-0.5}=1$. This rationale substantiates our selection of $\tau$ within the set $\{0.1,0.2,...,1\}$ as stated in Appendix D.2, aligning with the observed optimal performance in Figures 6(a)-(c). When addressing graphs with noisy structure, we may adjust the upper limit of $\alpha$ to $t\in(0,1)$, setting $\tau$’s maximum possible value to $\frac{t}{1-t}$. For graph benchmarking evaluations in this work, where extracting structural information is important, we practically set $t=0.5$. 2) For the non-local aggregation layer number $L$, a noticeable decline in model performance is observed when $L$ exceeds 10. This is attributed to the non-local nature of our new graph, which facilitates efficient information exchange between nodes. Exceeding a certain number of layers may potentially lead to oversmoothing, where there is an overemphasis on global information, thus degrading model performance. However, choosing the number of layers within a reasonable range generally ensures consistent and impressive model performance, as verified in Figures 6(j)-(l).
@@ 6.5. Parameter Study
Figure 5. Attention changing trends w.r.t. training epochs. Refer to Figure 9 in the Appendix E.2 for more visualizations.

### v5

@@ V-2 Non-local Spatial Aggregation
Once acquiring a suitable spectral filter $\hat{g}_{\psi}(\lambda)$, we compute the adapted new graph as $\hat{\mathbf{A}}^{\text{new}}=\mathbf{I}-\tau(\mathbf{U_{m}}g_{\psi}(\mathbf{\Lambda_{m}})^{-1}\mathbf{U_{m}}^{T}-\mathbf{I})$ by Eq. (6) where $\tau=\frac{\alpha}{1-\alpha}$ is a scaling parameter and a partial eigendecomposition [63] can be employed to produce a low-rank, robust structure for $\hat{\mathbf{A}}^{\text{new}}$ using only $m$ extremal eigenvalues. Equipped with this newfound graph, we proceed to perform non-local aggregation: $\mathbf{Z}^{(l)}=(1-\eta)\mathbf{Z}^{(0)}+\eta\hat{\mathbf{A}}^{\text{new}}\mathbf{Z}^{(l-1)}$, where $\eta$ refers to the update rate and $\mathbf{Z}^{(0)}=f_{\varphi}(\mathbf{X})$ and $l=1,2,\cdots,L$ denote layer number. The iteratively aggregated results are denoted as $\mathbf{Z}_{a}$. Recognizing the potential noise from the non-local nature of $\hat{\mathbf{A}}^{\text{new}}$, we apply a sparsification technique, leveraging a positive threshold $\epsilon$, and retain only essential elements outside the $\left[-\epsilon,\epsilon\right]$ interval. For clarity, this refined model is referred to as SAF-$\epsilon$.
V-3 Node-wise Prediction Amalgamation
@@
To leverage information from different graph domains, we employ an attention mechanism, allowing nodes to determine the importance of each space. This mechanism produces pairwise weights for a nuanced amalgamation during prediction. Specifically, the weight pair is computed as $\boldsymbol{\kappa}_{f}=\text{Sigmoid}(\mathcal{P}_{f}(\mathbf{Z}_{f}))$, $\boldsymbol{\kappa}_{a}=\text{Sigmoid}(\mathcal{P}_{a}(\mathbf{Z}_{a}))$ where $\boldsymbol{\kappa}_{f},\boldsymbol{\kappa}_{a}\in\mathbb{R}^{N}$ contain the weights for each node, and $\mathcal{P}_{f}(\cdot)$ and $\mathcal{P}_{a}(\cdot)$ are two different mappings from $\mathbb{R}^{C}$ to $\mathbb{R}$. For simplicity, we implement them using two one-layer MLPs. Given domain predictions $\mathbf{Y}_{f},\mathbf{Y}_{a}\in\mathbb{R}^{C}$, the final model prediction is attained as $\mathbf{Y}=\mathrm{diag}(\boldsymbol{\kappa}_{f})\cdot\mathbf{Y}_{f}+\mathrm{diag}(\boldsymbol{\kappa}_{a})\cdot\mathbf{Y}_{a}$ where a normalization $\left[\boldsymbol{\kappa}_{f},\boldsymbol{\kappa}_{a}\right]\leftarrow\frac{\left[\boldsymbol{\kappa}_{f},\boldsymbol{\kappa}_{a}\right]}{\max\{\|\{\boldsymbol{\kappa}_{f},\boldsymbol{\kappa}_{a}\}\|_{1},\delta\}}$ is performed beforehand to maintain $\boldsymbol{\kappa}_{f}+\boldsymbol{\kappa}_{a}=\mathbf{1}$ with small value $\delta$ preventing zero division. Similar schemes can be founded in works [64, 65, 66].
TABLE I: Statistics of real-world datasets. $F$ and $C$ denotes the number of features and classes. $\Delta$ represents graph diameter referring to the longest geodesic distance between nodes on the graph; for Penn94, due to multiple subgraphs, we report $\Delta$ of the largest connected component. We assess graph homophily using three metrics - edge homophily [46] $\mathcal{H}$, class homophily [67] $\mathcal{H}_{\text{class}}$, and adjusted homophily [68] $\mathcal{H}_{\text{adjusted}}$ which ranges from 0 (high heterophily) to 1 (high homophily).
[…]
Chameleon 2,227 36,101 2,325 5 11 0.23 0.06 0.03
Squirrel 5,201 217,073 2,089 5 10 0.22 0.03 0.01
Texas 183 309 1,703 5 8 0.11 0.00 -0.23
Cornell 183 295 1,703 5 8 0.30 0.05 -0.08
Actor 7,600 33,544 931 5 12 0.22 0.01 0.00
Cora 2,708 5,429 1,433 7 19 0.81 0.77 0.77
Citeseer 3,327 4,732 3,703 6 28 0.74 0.63 0.67
Pubmed 19,717 44,338 500 3 18 0.80 0.66 0.69
Minesweeper 10,000 39,402 7 2 99 0.68 0.01 0.01
Tolokers 11,758 519,000 10 2 11 0.59 0.18 0.09
Amazon-ratings 24,492 93,050 300 5 46 0.38 0.13 0.14
Roman-empire 22,662 32,927 300 18 6,824 0.05 0.02 -0.05
Penn94 41,554 1,362,229 5 2 8 0.47 0.05 0.02
V-4 Complexity Analysis
SAF augments spectral GNNs with non-local aggregation and node-wise amalgamation. The first part entails creating a new graph and information propagation. In SAF-$\epsilon$, these two steps are separated, culminating in $\mathcal{O}(N^{3}+N^{2}+nnz(\hat{\mathbf{A}}^{\text{new}})d)$ complexity, where $nnz$ denotes non-zero element count. Conversely, SAF, viewing non-local aggregation holistically, can reduce complexity to $\mathcal{O}(2dN^{2}+dN)$ when $d\ll N$. For node-wise amalgamation, its parallelizable nature ensures computational efficiency. Our method also requires eigendecomposition precomputation, which, though naively complex at $\mathcal{O}(N^{3})$, can be reduced to $\mathcal{O}(m^{2}+nnz(\hat{\mathbf{L}})m)$ using Lanczos method [63] with $m\ll N$ iterative steps. The results are also reusable for both training and inference. We present empirical studies on both time and space overheads in Section VI-J.
V-5 Practical Implications of Eigendecomposition
Eigendecomposition breaks down a matrix into its eigenvalues and eigenvectors, offering insights into matrix properties, especially for the graph Laplacian. Despite computational demands, this technique has attracted surging interest in the graph learning community due to its theoretical richness, and it can be practically expedited for larger graphs using Lanczos [63] and Sparse Generalized Eigenvalue [69] algorithms. Recent advancements also underscore its value in various applications such as graph positional encoding [70, 71, 72], spectral graph convolution [28], graph domain adaptation [73], graph robustness [74], graph expressivity [75, 76, 51, 27], and graph transformers [77, 78, 79, 43]. In line with these developments, our SAF also utilizes eigendecomposition to explicitly create a new graph, enabling efficient non-local aggregation with signed weights to tackle long-range dependency and graph heterophily.
@@ V-5 Practical Implications of Eigendecomposition
TABLE II: Semi-supervised node classification accuracy (%) $\pm$ 95% confidence interval.
@@
Improv. [^3] Improv. indicates the relative improvement of our SAF over its base model, BernNet [13]. For alternative implementation using ChebNetII [15] as backbone, please refer to Section VI-H. 14.56% 9.68% 15.37% 13.99% 3.97% 1.62% 1.86% 0.68%
VI Experiments
VI-A Datasets and Experimental Setup
VI-A1 Datasets
We conduct experiments on 13 real-world datasets with detailed statistics provided in Table I. In certain compact sections of this paper, we use four-letter abbreviations for dataset names.
  • Chameleon and Squirrel are Wikipedia networks collected by [80] where nodes are web pages connected by mutual links. We utilize the labels from [39], dividing nodes into five classes based on their average monthly traffic.
  • Texas and Cornell are webpage datasets from WebKB project, containing nodes from five classes (student, project, course, staff, and faculty) and connected by hyperlinks. This study uses the preprocessed version by [39].
  • Actor is an actor co-occurrence network induced from the film-director-actor-writer network [81] by [39]. In this network, nodes are actors divided into five classes, and edges represent actor co-occurrence on Wikipedia pages.
  • Cora, Citeseer, and Pubmed are citation network benchmarks [82, 83] with nodes as scientific papers and edges as undirected citations. Each node is assigned with a class label by research topic and bag-of-word features.
  • Minesweeper, Tolokers, Amazon-ratings, and Roman-empire are recently proposed benchmarks by [84] for specifically evaluating GNN models under heterophily.
  • Penn94 is a large-scale friendship network [85] from the Facebook 100 networks, where nodes denote students labeled by reported genders and posses features such as major, second major/minor, dorm/house, year, and high school. This work uses the version preprocessed by [67].
VI-A2 Baselines
We compare SAF with 21 models: (1) MLP; (2) Basic GNNs: GCN [10] and APPNP [37]; (3) Spectral GNNs: ARMA [36], GPR-GNN [12], BernNet [13], ChebNetII [15], JacobiConv [16], Specformer [43], LON-GNN [18], OptBasisGNN [19] and FLODE [44]; (4) Spatial GNNs: GCNII [25], PDE-GCN [38], NodeFormer [41], GloGNN++ [56] and LRGNN [86]; (5) Unified GNNs: GNN-LF [24], GNN-HF [24], ADA-UGNN [23] and FE-GNN [27].
VI-A3 Implementation Details
To follow [13, 15, 16], we fix $K=10$. For each dataset, we perform a grid search to tune the hyper-parameters of all models. With the best hyper-parameters, we train models with Adam optimizer [87] in 1,000 epochs using early-stopping strategy and a patience of 200 epochs, and report the mean classification accuracies with a 95% confidence interval on 10 random data splits. As [15] have made a thorough evaluation and share the same experimental protocol with us, we leverage their results for models: MLP, GCN, APPNP, ARMA, GPR-GNN, BernNet, ChebNetII, GCNII and PDE-GCN on datasets – Chameleon, Squirrel, Texas, Cornell, Actor, Cora, Citeseer, and Pubmed. For JacobiConv, LON-GNN, and OptBasisGNN, we also report the results from their papers [16, 18, 19]. Our codes will be made available if the paper could be accepted.
VI-A4 Hyper-parameters Setting
[…]
TABLE III: Full-supervised node classification accuracy (%) $\pm$ 95% confidence interval.
@@ VI-A4 Hyper-parameters Setting
Improv. [^3] Improv. indicates the relative improvement of our SAF over its base model, BernNet [13]. For alternative implementation using ChebNetII [15] as backbone, please refer to Section VI-H. 6.77% 12.61% 2.13% 1.15% 1.27% 1.36% 1.37% 3.01%
VI-B Semi-supervised Node Classification.
@@
In this task, we follow the experimental protocol established by [15] and compare our models with MLP, two basic GNNs, eight popular polynomial spectral GNNs, and five unified GNNs. For data splitting on homophilic graphs (Cora, Citeseer, and Pubmed), we apply the standard division [89] with 20 nodes per class for training, 500 nodes for validation, and 1,000 nodes for testing. On the other five heterophilic graphs, we leverage the sparse splitting [12] with 2.5%/2.5%/95% samples respectively for training/validation/testing. The results are reported in Table II, where the best results are bold and the underlined letters denote the second highest accuracy. We first observe that both SAF and SAF-$\epsilon$ substantially boosts its base model, BernNet, with gains reaching a notable 15.37%. This impressive enhancement is credited to their capacity to effectively exploit the task-beneficial information, which is implicitly encoded by spectral filtering in the spatial domain. This ability is particularly advantageous in contexts with limited supervision, where it allows effective leveraging of extra prior knowledge during training. Generally, our models outperform competitors on all datasets except for Chameleon and Squirrel, where SAF maintains a second-place rank with considerable improvements on BernNet by 14.56% and 9.68%. In these cases, ChebNetII initially surpasses our model, yet, with more training samples, our SAF manages to beats it by margins of 3.93% and 6.28% (see Table III). Moreover, SAF-$\epsilon$ averagely delivers better results than SAF by using thresholding sparsity to reduce non-local noise. However, this enhancement also incurs higher computational costs, as illustrated in both Section V-4 and Section VI-J.
VI-C Full-supervised Node Classification.
@@
To bolster our evaluation, we expand the previously compared baselines to include five cutting-edge spatial GNN models: GCNII & PDE-GCN capturing long-range dependency, and NodeFormer, GloGNN++ & LRGNN, which go further by not only capturing long-range dependencies but also effectively addressing graph heterophily. For all datasets, we randomly divide them into 60%/20%/20% for training/validation/testing by following [13, 15]. Table III summarizes the mean classification accuracies. Our methods demonstrate superior performance across most datasets, with an exception on Squirrel where they achieve comparable results to Specformer. This notable performance is primarily attributed to our SAF’s effective non-local aggregation, utilizing signed edge weights to model global label relationships. This enables our methods to outperform GNNs that are specifically tailored for long-range dependency and/or graph heterophily.
@@
Figure 4: Attention changing trends w.r.t. training epochs.
VI-D Analysis of Attention Trends
We analyze the changing trends of the pair-wise attention weights during training SAF on Squirrel and Cora datasets. From Fig. 4(b-c), the average weights for filtering and aggregation start similarly but diverge throughout training, showing different trends in heterophilic and homophilic graphs. On the heterophilic graph Squirrel, both weights converge to similar values, demonstrating their mutual importance in modeling complex connectivity. Conversely, $\boldsymbol{\kappa}_{f}$ becomes dominant on the homophilic graph Cora due to the sufficiency of node proximity information for label prediction, thereby diminishing the relevance of $\boldsymbol{\kappa}_{a}$ and non-local aggregation.
Figure 5: Ablation study of SAF framework on five datasets.